OpenStudy (anonymous):

The solution of quantity of x plus 5 divided by 2 equals 15 is 25.

OpenStudy (anonymous):

x + 5 ------ = 15 2 solve for x: x = 15 x 2 - 5 x = 30 - 5 x = 25